the rate equation for the second order reaction of E and S is

sapling learning The rate equation for the second-order reaction of E and S is where kcat/KM is the second-order rate constant, which can be used to compare catalytic efficiencies cat What is the diffusion-controlled limit in aqueous solution? O 107 to 108 M s O 108 to 10 M-1s-1 O 10-6 to 10-5 M O 109 to 1010 M s1 O 108 to 10° MIs O 10-7 to 10-8 M s-1 O 10-9 to 10-8 M-s1what os the diffusion controlled limit in aqueous solution
Answer #1
Concepts and reason

Enzyme and substrate collides with each other in enzyme catalyzed reactions. The frequency of enzyme-substrate collision is in the range of 108to109M1s1{10^8}{\rm{ to 1}}{{\rm{0}}^9}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}} for a second order enzyme catalyzed reaction.

The ratio of kcat/KM{k_{{\rm{cat}}}}/{K_{\rm{M}}} cannot be greater than 108to109M1s1{10^8}{\rm{ to 1}}{{\rm{0}}^9}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}} for a second order reaction to formES{\rm{ES}}, which is diffusion-controlled limit.

Fundamentals

For a second order reaction between an enzyme (E) and substrate (S) with following mechanism:

Initialvelocity,V0=[S]KM+[S]Vmax{\rm{Initial velocity, }}{V_{\rm{0}}} = \frac{{[{\rm{S}}]}}{{{K_{\rm{M}}} + [{\rm{S}}]}}{V_{\max }}

Here, Vmax=kcat[E]total{V_{\max }} = {k_{{\rm{cat}}}}{[{\rm{E}}]_{{\rm{total}}}} and [E]total=[E]+[ES]{[{\rm{E}}]_{{\rm{total}}}} = [{\rm{E}}] + [{\rm{ES}}] . Michaelis-Menton constantKM=k2+k1k1{K_{\rm{M}}} = \frac{{{k_2} + {k_{ - 1}}}}{{{k_1}}}When [S]<<KM[{\rm{S}}] < < {K_{\rm{M}}} then KM+SKM{K_{\rm{M}}} + {\rm{S }} \approx {K_{\rm{M}}} . The expression for V0{V_{\rm{0}}} becomes as follows:

V0=[S]KM+[S]Vmax=kcatKM[E]total[S]\begin{array}{c}\\{V_{\rm{0}}} = \frac{{[{\rm{S}}]}}{{{K_{\rm{M}}} + [{\rm{S}}]}}{V_{\max }}\\\\ = \frac{{{k_{{\rm{cat}}}}}}{{{K_{\rm{M}}}}}{[{\rm{E}}]_{{\rm{total}}}}[{\rm{S}}]\\\end{array}

The ratio kcat/KM{k_{{\rm{cat}}}}/{K_{\rm{M}}} is the second-order rate constant for the formation of ES{\rm{ES}}.

Following are second rate constants in the given set of rate constants.

108to109M1s1109to108M1s1\begin{array}{l}\\{10^8}{\rm{ to 1}}{{\rm{0}}^9}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}}\\\\{10^{ - 9}}{\rm{ to 1}}{{\rm{0}}^{ - 8}}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}}\\\end{array}

The diffusion-controlled limit in aqueous solution is 108to109M1s1{10^8}{\rm{ to 1}}{{\rm{0}}^9}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}}.

Ans:

The diffusion-controlled limit in aqueous solution is 108to109M1s1{10^8}{\rm{ to 1}}{{\rm{0}}^9}{\rm{ }}{{\rm{M}}^{ - 1}}{{\rm{s}}^{ - 1}}.